Power Up has created an indestructible charger cord that sells for $20 per unit. The companys contribution format income statement for the most recent month is given below Sales (24,400 units at s$20.00 per unit Variable expenses Contribution margin Fixed expenses Net operating income $488,000 146,400 $341,600 193,000 $148,600 1. What is the companys contribution margin per unit? (1 Point) Submit Answer Tries 0/10 2. What is Power Ups break-even amount in sales dollars? (1 Point) Submit Answer Tries 0/10 3. The sales manager estimates that sales will increase $35,000 during the coming year due to increased demand. If the sales manager is right, by how much should net operating income increase? Present an increase as a positive number and a decrease in NOI as a negative number. (1 Point) Submit Answer Tries 0/10 4. Compute the degree of operating leverage at the original level of sales. COMPUTE TO FOUR DECIMAL PLACES.(1 Point) Submit Answer Tries 0/10 5. The president of the company expects sales to increase by 10% next year. By how much should net operating income increase (in dollars)? Hint: Use the original data (1 Point) Submit Answer Tries o/10
Answer #1
1
Contribution margin per unit = 341600/24400= $14
2
CM ratio =341600/488000= 70%
Break-even amount in sales dollars = Fixed expenses/CM ratio = 193000/70%= $275714 or $275714.3
3
Net operating income increase = 35000*70%= $24500
4
Degree of Operating leverage = Contribution margin/Net operating income = 341600/148600= 2.2988
5
Net operating income increase = 488000*10%*70%= $34160